You’re prepared to make monthly payments of $340, beginning at the end of this month, into an account that pays 6 percent interest compounded monthly. How many payments will you have made when your account balance reaches $20,000?
1
Expert's answer
2020-07-01T18:57:37-0400

A = P ( 1+ r/n )12t


Compounded monthly:

1 year = 12 months

= 0.06/12


= 0.005%


20000 = 340 ( 1+ 0.06/12)12t


2000/340 = ( 1+ 0.005)12t


5.8824 = 1.00512t


Log 5.8824 = Log 1.00512t


Log 5.8824 / Log 1.005 =


12t ( Log 1.005/ Log 1.005)


355.2782 = 12t


t = 355.2782/12 = ~ 29.6065months


= 29 months 18 days